how to purchase a cover for a swimming pool that has a 3-foot walkway all the way around it. The cover must fit over the pool and the walkway.

Answer by CharlesG2(834)

how to purchase a cover for a swimming pool that has a 3-foot walkway all the way around it. The cover must fit over the pool and the walkway.
x = length of pool
y = width of pool
x + 3 = length of pool + walkway
y + 3 = width of pool + walkway
A area of pool + walkway = (x + 3)(y + 3) = xy + 3x + 3y + 9 square feet
find out cost per square foot
take cost per square foot * area = cost for cover